OpenStudy (anonymous):

A sporting goods store has added a section with trail foods and snacks. Raisins cost $5.00 per pound and peanuts cost $3.80 per pound. Raisins and peanuts are going to be sold in a mixture of 30 pounds which will cost $4.05 per pound? How many pounds of raisins will be in the mix?

OpenStudy (anonymous):

Answer: 23.75 lbs. Let p = lbs of peanuts in mixture r = lbs of raisins in mixture Solve the following system: p + r = 30 5p + 3.8r = 121.5 (4.05 x 30) p = 6.25 r = 23.75